### v4

# Enhancing Edge Detection by Texture Handling Architecture and Noiseless Training Data

Abstract
Image edge detection (ED) is a fundamental task in computer vision. While convolution-based models have significantly advanced ED performance, achieving high precision under strict error tolerance constraints remains challenging. Furthermore, the reliance on noisy, human-annotated training data limits model performance, even when the inputs are edge maps themselves. In this paper, we address these challenges in two key aspects. First, we propose a novel ED model incorporating Cascaded Skipping Density Blocks (CSDB) to enhance precision and robustness. Our model achieves state-of-the-art (SOTA) performance across multiple datasets, with substantial improvements in average precision (AP), as demonstrated by extensive experiments. Second, we introduce a novel data augmentation strategy that enables the integration of noiseless annotations during training, improving model performance, particularly when processing edge maps directly. Our findings contribute to a more precise ED architecture and the first method for integrating noiseless training data into ED tasks, offering potential directions for improving ED models. Codes can be found on https://github.com/Hao-B-Shu/SDPED.
keywords:
Edge Detection, Convolutional Neural Network, Data Augmentation, Noiseless Data, Evaluation Benchmark.
[…]
organization=Shenzhen University,city=Shenzhen, postcode=518060, state=Guangdong, country=P. R. China
{highlights}
An edge detection model based on CSDB.
A data augmentation method allowing noiseless data.
A discussion of the choice of evaluation standard.
Significant improvements over high standard evaluation.
Extensive experiments implementation.
1 Introduction
Image edge detection (ED) is a fundamental task in computer vision and has been studied for decades, with extensive applications in high-level tasks such as object detection[1] and image segmentation[2]. Early ED methods, such as the Sobel operator[3] and the Canny algorithm[4], relied on gradient-based features. Later, more sophisticated statistical approaches were introduced[5; 6]. In recent years, machine learning has become indispensable in nearly all computer vision tasks, leading to the emergence of numerous learning-based ED algorithms[7; 8; 9; 10; 11; 12]. Currently, convolutional neural network (CNN)-based models achieve state-of-the-art (SOTA) performance in ED tasks [13; 14; 15; 16; 17; 18; 19], while recently, transformer-based approaches have also been explored[20].
The goal of ED is to identify edges in an image, which typically correspond to pixels with significant changes in intensity compared with their neighbors. However, textural areas can also contain rapid intensity changes that do not correspond to edges. Classical ED algorithms such as Canny[4], despite their efficiency, often fail to differentiate between edges and textures due to their inability to adapt to varying data distributions. Consequently, learning-based models, particularly CNNs, have become the dominant approach. Over the past decade, CNN-based ED models have demonstrated remarkable progress, with models such as HED[14], RCF[21], RINDNet[22], BDCN[23], DexiNed[19], and even lightweight models such as TEED[24] achieving $F$-scores comparable to or exceeding human performance in standard evaluation procedures.
Despite these advancements, several challenges remain in current ED research. In this paper, we focus on the following two key issues:
  1. 1.
    Limited precision of existing models: Current ED models achieve high scores only when evaluated with a large error tolerance. For example, the standard error tolerance radius is approximately 4.3 pixels for BSDS500[6], 11.1 pixels for BIPED[25], and even 16.3 pixels for MDBD[26]. However, when stricter evaluation criteria are applied, the performance of these models drops significantly. Hence, there is a pressing need for more precise ED models that perform well under stricter evaluation settings.
  2. 2.
    Suboptimal predictions on noiseless inputs: Existing models often produce suboptimal edge maps even when given noiseless input data, namely ground-truths themselves [^1] In this paper, we define noiseless data as ground-truth edge maps, while noisy data refers to standard images containing textures. Models should ideally perform better on noiseless inputs than on noisy ones., as shown in Fig 1. Ideally, an ED model should yield nearly perfect predictions on noiseless inputs since they are already well-defined edges. Yet, most models generate results on noiseless inputs that are similar to predictions on textured images. Addressing this issue is critical for improving ED model reliability.
Figure 1:  Predictions of three previous ED models on the BRIND dataset. $Off_{E}$ denotes models retrained using the official data augmentation and tested on edge maps, while $Our_{E}$ represents models retrained using our proposed data augmentation (detailed in Section III) and tested on edge maps. Similarly, $Off_{O}$ and $Our_{O}$ correspond to models trained with respective augmentation strategies but tested on standard images. Our augmentation method produces sharper predictions on edge maps and potentially enhances performance on standard images.
To address these challenges, we propose a novel ED model incorporating cascaded skipping density blocks (CSDB). Density blocks (DBs) have demonstrated effectiveness in tasks such as image super-resolution [27], however, they perform suboptimally when directly applied to ED. Therefore, we introduce CSDB, a structural refinement tailored for ED tasks, and design new ED models based on this architecture. Additionally, to improve model performance on noiseless data, we introduce a novel data augmentation strategy that incorporates noiseless samples. Unlike conventional augmentation methods that rely solely on textured images with noisy labels, our approach allows models to learn directly from noiseless data, thereby mitigating the issue discussed in 2) and Fig. 1.
The key contributions of this work are as follows:
  • 1.
    Development of high-precision ED models: We introduce a novel ED model, Skipping Density Precise Edge Detection (SDPED), which achieves SOTA performance with fewer parameters. SDPED employs deep architectures with CSDB, a structure previously unexplored in ED. Notably, our models do not require multi-scale supervision, challenging the widely held belief that such supervision is essential for ED.
  • 2.
    Introduction of a novel data augmentation strategy: We propose a data augmentation method that incorporates noiseless edge maps into training, enabling models to achieve improved performance on noiseless data. This method provides a potential solution to the issue described in 2) and Fig. 1.
  • 3.
    Comprehensive experimental validation: We conduct extensive experiments to validate our approaches, demonstrating their effectiveness through quantitative and qualitative analyses.
2 Review of Previous Work
This section reviews previous works in the field of ED, including datasets, pre-processing techniques, CNN-based ED models, loss functions, and evaluation metrics.
2.1 Datasets, Label Refinement, and Data Augmentation
Early ED datasets were primarily derived from segmentation tasks, such as BSDS300 and its later extension, BSDS500[28], which remains one of the most widely used benchmarks. BSDS500 consists of 500 RGB images annotated by multiple users and is frequently utilized for both segmentation and boundary detection tasks. Another well-known dataset is NYUD/NYUD2[29], which contains hundreds of thousands of RGB-D images, originally designed for indoor scene analysis. Among these, 1,449 images are annotated with segmentation categories. Other commonly used datasets include PASCAL VOC[30], Microsoft COCO[31], SceneParse150[32], and Cityscape[33]. However, since these datasets were not explicitly designed for ED, their labels may lack precision in edge annotations.
With the increasing focus on ED tasks, several dedicated datasets have been proposed. In 2016, the MDBD dataset[26] was introduced, containing 100 high-resolution (1280$\times$720) images, each annotated by multiple annotators. While initially designed for boundary detection, it has been widely adopted in ED research. In 2020, the BIPED dataset[25] was specifically developed for ED tasks, featuring 250 high-resolution (1280$\times$720) RGB images with precise edge annotations, later refined in its BIPED2 version. In 2021, BRIND[22] was introduced as a re-annotation of BSDS500, categorizing edges into four types: reflectance, illumination, normal, and depth edges. More recently, UDED[24], proposed in 2023, consists of 29 high-quality images with carefully labeled edges, curated from multiple datasets.
A fundamental challenge in all ED datasets is annotation noise. Since edge labels are based on human annotations, perfect accuracy is unattainable, and inconsistencies arise both between annotators and even within the same annotator across multiple attempts. As a result, various studies have focused on refining human-labeled data or enhancing model robustness to noisy annotations[34; 35].
Regardless of the dataset used, data augmentation is widely recognized as essential in ED tasks. Common augmentation techniques include rotation, flipping, cropping, resizing, and gamma transformations. However, augmentation strategies vary across studies. For instance, in HED[14], images are rotated at 16 different angles, flipped, and cropped to obtain 32 augmented images per sample. In RIND[22], only four rotations with flips (without cropping) are applied, resulting in eight augmented images per sample. In contrast, Dexi[19] employs a more aggressive augmentation strategy, incorporating 16 rotations, flips, cropping, and three gamma transformations, generating 288 augmented images per sample. While extensive augmentation increases data diversity, its impact on model performance remains an open question, but excessive augmentation can significantly increase training costs.
2.2 CNN-based ED Models and Loss Functions
Currently, CNN-based methods dominate ED tasks. The HED model[14] introduced a multi-scale and multi-level output strategy, where predictions at different levels are fused, and the loss is computed as the average of all supervised outputs. RCF[21] extends this approach by integrating multi-scale outputs with feature fusion across different layers, generating side outputs to enhance edge representation. BDCN[23] applies layer-specific supervision, further improving ED accuracy. PiDiNet[17] introduces pixel-difference convolutions, while RIND[22] predicts four edge types through a three-stage processing pipeline. Dexi[19] leverages extremely dense skip connections to enhance feature propagation.
Most ED models struggle to produce crisp edges without additional post-processing techniques such as non-maximal suppression (NMS). This issue may arise from the use of down-sampling operations such as max-pooling, as well as the widespread use of weighted binary cross-entropy (WBCE) as the loss function. Various strategies have been proposed to address this challenge. For instance, [36] suggests improving edge sharpness by aggregating features at larger scales during up-sampling, while [37] explores the effectiveness of Dice loss in ED tasks. More recently, tracing loss[18] has emerged as a promising alternative and has been adopted in models such as TEED[24]. However, these solutions often introduce trade-offs, requiring careful evaluation in different scenarios.
2.3 Evaluation Metrics for ED
The standard evaluation framework for ED models includes three key metrics: Optimal Dataset Score (ODS), Optimal Image Score (OIS), and Average Precision (AP). These metrics are derived from the $F_{\beta}$-score, defined as:
$F_{\beta}=\frac{(1+\beta^{2})\times\text{Precision}\times\text{Recall}}{\beta^{2}\times\text{Precision}+\text{Recall}},$ (1)
where Precision is calculated as:
$\text{Precision}=\frac{TP}{TP+FP},$ (2)
and Recall is given by:
$\text{Recall}=\frac{TP}{TP+FN}.$ (3)
Here, TP (true positives) refers to correctly predicted edge pixels, FP (false positives) refers to non-edge pixels incorrectly classified as edges, and FN (false negatives) refers to missed edge pixels. The parameter $\beta$, typically set to 1, balances precision and recall.
Most ED models utilize the evaluation algorithm from[7], which includes a hyperparameter, maxDist, representing the maximum allowable error tolerance distance. The error tolerance distance is expressed as a fraction of the diagonal length of the image. For example, in BSDS500, where images have a resolution of 321$\times$481 pixels, setting $\textit{maxDist}=0.0075$ results in an error tolerance of approximately 4.3 pixels, computed as $\sqrt{481^{2}+321^{2}}\times 0.0075$. Notably, the choice of maxDist is often inconsistent across datasets. For instance, in BSDS500[28], maxDist is set to 0.0075 (about 4.3 pixels), whereas in BIPED[19], the same value translates to 11.1 pixels due to higher image resolution. In MDBD[26], maxDist is often increased to 0.011, allowing a tolerance of approximately 16.3 pixels. A stricter evaluation standard would involve reducing the error tolerance distance, though such adjustments require additional discussions.
3 Methodology
This section presents the core methodology of our approaches, including the architecture of the proposed SDPED model and an improved data augmentation strategy designed to enhance model performance on noiseless data. First, we introduce the SDPED model, which leverages the CSDB structure to improve edge feature extraction. Then, we discuss our modified data augmentation method, which incorporates noiseless data into training to mitigate the limitations caused by noisy human annotations.
3.1 The SDPED Model
The performance of ED models largely depends on their ability to suppress textural regions while preserving true edges. Since textures lack a strict definition, designing theoretically sound architectures to handle textures remains a challenge. However, ED research has progressed more slowly compared to other computer vision tasks, such as image super-resolution (SR), and the latter has produced numerous architectures experimentally validated for texture processing. Borrowing structural designs from related tasks may, therefore, be a promising direction for improving ED models.
Based on this observation, we propose the SDPED model, illustrated in Fig. 2, with additional visualizations in Fig. LABEL:fig:Legend and 3. The model employs CSDBs in the main processing path, generating multiple outputs that are subsequently fused by a dedicated fusing block. The Dense Block (DB) structure has been widely used in SR tasks [38; 27], demonstrating strong performance in texture processing. Since texture-rich regions pose challenges in both SR and ED tasks, it is reasonable to explore similar architectures for ED. However, to better suit ED-specific requirements, we introduce modifications by replacing DB with CSDB, which is better tailored for this application.
Additionally, most existing ED models allocate tens of millions of parameters to feature extraction but rely on a single $1\times 1$ convolution layer for feature fusion. This imbalance may lead to underutilization of extracted features, potentially limiting performance. To address this, unlike previous models that typically fuse multiple outputs using a single $1\times 1$ convolution layer, the proposed model employs a fusing block consisting of three convolution layers, including a $3\times 3$ convolution layer, to enhance feature integration.
Figure 2: Structure of the SDPED Model: The blue square represents the $3\times 3$ convolutional layer, the purple square represents the $1\times 1$ convolutional layer, the pink square represents the Leaky Relu, the black square represents the Sigmoid, $\oplus$ represents pixel-wise addition, and represents the concatenation. The top path represents the main processing pipeline, where the first two convolution layers output 32 and 64 feature maps, respectively. The last two convolution layers maintain 64 input and output features. The model comprises $n+2$ layers, including the feature extractor block, each CSDB unit, and the final processing block, all of which produce intermediate outputs fused into 21 features. These intermediate outputs are concatenated, resulting in a total of $21\times(n+2)$ features for the final fusing layer. The bottom path represents the fusing block, where the concatenated features pass through a $3\times 3$ convolution layer followed by two $1\times 1$ convolution layers. The output feature dimensions of these layers are 256, 512, and 1, respectively.
The model is trained using the weighted binary cross-entropy (WBCE) loss function, defined as:
$L_{WBCE}(\hat{Y},Y)=-\alpha\sum_{y_{i}\in Y^{+}}\log(\hat{y_{i}})-\lambda(1-\alpha)\sum_{y_{i}\in Y^{-}}\log(1-\hat{y_{i}}),$ (4)
where $\hat{Y}$ represents the model’s prediction, $Y^{+}$ is the set of positive samples (edge pixels in the ground-truth), and $Y^{-}$ is the set of negative samples (non-edge pixels in the ground-truth). The weighting factor $\alpha$ is computed as:
$\alpha=\frac{|Y^{-}|}{|Y|}.$ (5)
Figure 3: Cascading Skipping Dense Block (CSDB): Each CSDB block consists of three cascaded SDB blocks connected via skip connections. Within a single SDB block, there are five convolution layers. The first four layers output 32 feature maps, while the final layer produces 64 feature maps. The input feature dimensions grow progressively from 64 to 96, 128, 160, and 192 across the first four layers.
3.2 Enhancing Performance on Noiseless Data
Existing models, along with their data augmentation techniques, often fail to yield satisfactory results even on noiseless data. Edge maps inherently contain higher-quality edge information than raw images, and thus predictions on such noiseless data should ideally be sharper and more precise. However, as shown in Fig. 1, previous models tend to produce similar outputs for both noisy and noiseless inputs.
The primary reason for this limitation is that existing models are trained on datasets where label noise is inevitable. Like all supervised learning methods, ED models are highly dependent on the quality of training data. However, obtaining high-quality ED datasets is particularly challenging, as edge annotations are manually labeled and inherently noisy, often containing missing or erroneous labels. This labeling noise can significantly degrade model performance, even when evaluated on high-quality test data.
The most straightforward solution to this problem would be incorporating noiseless data into the training set. However, since human annotations are currently the only feasible method for generating ground-truth labels, obtaining completely noiseless data appears impractical.
Nevertheless, a potential solution arises by directly using ground-truth labels as training samples. ED labels are inherently noiseless when treated as input images with their own ground-truth since the edge map of an edge map—regardless of how it is generated—remains unchanged. Leveraging this property, we integrate noiseless data into the training process by using labels as input images, effectively enhancing dataset quality without requiring additional annotations. With this modified data augmentation strategy, models can generate more accurate and visually appealing predictions, particularly on noiseless inputs, as demonstrated in Fig. 1.
4 Experiment
This section is dedicated to experiments. We begin by outlining the dataset settings and training methods, followed by a discussion of the evaluation metrics. Finally, we provide the experimental results and ablation studies.
4.1 Data and Training Setting
This subsection presents the dataset settings and training methods.
4.1.1 Dataset
Our experiments were conducted on four datasets: BRIND[22], UDED[24], MDBD[26], and BIPED[25]. Since BRIND provides multiple annotations for each image, we consolidate all labels into a single binary edge map [^2] We merge all available ground-truths into a unified binary map, marking a pixel as an edge if it is labeled as such in any of the annotations.. For MDBD, we use annotation 3 as the ground-truth [^3] Annotation 3 is chosen arbitrarily; alternative combinations of the six available annotations were tested, but the ground-truth quality was found to be unsatisfactory.. UDED and BIPED each provide a single label per image, which is used as the ground-truth. Notably, among these datasets, UDED exhibits the highest annotation quality, while BRIND offers the most diverse image collection.
4.1.2 Data Augmentation
To enhance data diversity, training images are recursively split in half until their height and width are both below 640 pixels. Each resulting sub-image undergoes eight augmentation transformations: rotations of 0o, 90o, 180o, and 270o, combined with horizontal flips. Furthermore, to improve performance on noiseless data, the original labels themselves are added to the training set, with their ground-truths set to their own values.
4.1.3 Training Detail
All models are trained using the Adam optimizer with a weight decay of $10^{-8}$. The initial learning rate is set to $10^{-4}$ and is reduced by a factor of 10 every 50 epochs. The batch size is fixed at 8, and the balance parameter in the loss function is set to $\lambda=1.1$ (please refer to Equation (4)). During training, images are randomly cropped to 320$\times$320 pixels, and the dataset is refreshed every five epochs. To avoid resizing operations during training, the two UDED images with widths below 320 pixels are designated as test images. The number of training epochs is set as follows: BRIND: 100 epochs, UDED: 150 epochs, MDBD: 100 epochs, BIPED: 75 epochs.
4.2 Benchmark
4.2.1 Evaluation benchmark
To evaluate model performance, we adopt the widely used methodology from [7], which computes ODS, OIS, and AP under an error tolerance distance. However, previous studies have employed different error tolerance distances: approximately 4.3 pixels for BRIND [28], 11.1 pixels for BIPED [19], and up to 16.3 pixels for MDBD [26]. Many researchers retain the default evaluation setting, where the error tolerance distance is a fixed ratio of the image’s diagonal length, inadvertently leading to inconsistent benchmarks [^4] For instance, in BRIND (image resolution: $321\times 481$), an error tolerance ratio of 0.0075 corresponds to approximately 4.3 pixels, whereas in BIPED (resolution: $720\times 1280$), the same ratio allows for 11.1 pixels..
Using a diagonal-length ratio as the error tolerance criterion is problematic because it relaxes the evaluation standard as image resolution increases. For example, in the BSDS500 dataset (resolution: 321$\times$481), a ratio of 0.0075 results in a 4.3-pixel tolerance, whereas concatenating four such images into a single 642$\times$962 image effectively doubles the tolerance to 8.6 pixels. Consequently, artificially enlarging images can inflate evaluation scores without improving actual ED performance. To ensure a resolution-independent evaluation, we propose setting the error tolerance distance as a fixed pixel value rather than a proportional ratio.
Moreover, the choice of an appropriate error tolerance threshold for edge detection tasks has been largely overlooked. We argue that models should be evaluated using a low error tolerance distance, since poor performance under a strict threshold but high scores under a relaxed one indicate that erroneous predictions contribute disproportionately to the final score [^5] For example, a prediction missing a single edge pixel should score higher than one missing the same pixel while introducing an additional false edge elsewhere. However, a loose error tolerance threshold can increase the likelihood of the latter receiving a higher score, as the erroneous edge pixel may still fall within the tolerance range..
However, since edge annotations in current datasets are manually labeled, they often exhibit a one-pixel deviation due to human annotation inconsistencies. As illustrated in Fig. 4, edges are typically marked as one pixel wide, even though true edges should be two pixels wide. This means that ground-truth labels may vary by up to one pixel horizontally, vertically, or diagonally. Therefore, while a zero-pixel error tolerance represents the strictest possible standard, a one-pixel tolerance is more realistic given the labeling inconsistencies, whereas allowing more than one pixel of tolerance risks overestimating model accuracy. Based on these considerations, we propose using a one-pixel threshold for rigorous evaluation.
Figure 4: Edge annotation inconsistencies: (a) A typical image region; (b) and (c) illustrate different annotation styles, where black squares denote edge pixels. In (b) and (c), non-corner edge pixels may deviate by one pixel horizontally or vertically, while corner edge pixels may shift diagonally by one pixel.
4.2.2 Evaluation Settings
All models are evaluated on full-resolution test images [^6] Unless explicitly stated, noiseless data are excluded from testing.. The standard evaluation algorithm from [7] is used, with the strictest error tolerance setting: one pixel in the diagonal direction (approximately 1.42–1.99 pixels in horizontal and vertical directions). The corresponding error tolerance distances are set as follows: BRIND: 0.003, MDBD and BIPED: 0.001, UDED (which contains images of varying resolutions): 0.004, ensuring a minimum of one-pixel tolerance for the smallest image ($250\times 361$ resolution).
As noted in [39], many models achieve SOTA performance on a specific dataset at the cost of generalization, leading to significant performance degradation when tested on different datasets [^7] This issue can arise even when only the training-test partitioning of the same dataset is altered.. To ensure a comprehensive and reliable evaluation, we conduct experiments on multiple datasets with various dataset partitions. We use the notation ’Dataset-$P_{k}$-a-b-Ec’ to index dataset partitions, where: Dataset refers to the dataset name, $P_{k}$ denotes the $k$-th partition, $a$ and $b$ represent the number of training and test images, respectively, and $c$ indicates the number of training epochs. For example, ’BRIND-$P_{1}$-250-250-E100’ refers to experiments conducted on the BRIND dataset using the first partition, where 250 images are used for training, 250 for testing, and the model is trained for 100 epochs.
4.2.3 Comparison Settings
The SDPED model is compared against several methods: HED[14], RCF[21], RIND [22], BDCN[23], and Dexi[19]. Since the SDPED architecture can be configured with different numbers of CSDB blocks, we denote a model with $n$ CSDB blocks as SDPED_$\_n$, with SDPED_$\_7$ being the standard for comparison. Moreover, for fairness, all models are trained under identical conditions, including the same dataset, optimizer, normalization strategy, and prediction-saving procedure.
Table 1:   Results on BRIND with error toleration 0.003 (1 pixel on the diagonal): The best results are highlighted in blue, while the second-best results are marked in red. In all three experiments and across all three benchmarks, our model (SDPED_$\_7$) achieves the best results. The improvements compared to the second-best results (for each partition and benchmark, regardless of the previous model) are also listed. On average, our model improves scores by approximately $2.7\%$ in ODS, $3.1\%$ in OIS, and $6.9\%$ in AP.
Partitions BRIND-$P_{1}$-250-250-E100 BRIND-$P_{2}$-300-200-E100 BRIND-$P_{3}$-400-100-E100
Benchmarks ODS OIS AP ODS OIS AP ODS OIS AP
HED (2015) 0.639 0.651 0.575 0.657 0.669 0.597 0.663 0.675 0.605
RCF (2017) 0.643 0.653 0.573 0.663 0.672 0.598 0.665 0.674 0.603
RIND (2021) 0.624 0.634 0.509 0.651 0.661 0.610 0.649 0.662 0.590
BDCN (2022) 0.639 0.652 0.567 0.645 0.659 0.576 0.662 0.673 0.587
Dexi (2023) 0.639 0.648 0.548 0.652 0.663 0.572 0.663 0.674 0.585
SDPED_$\_7$(Ours) 0.660 0.672 0.613 0.677 0.693 0.643 0.687 0.698 0.658
[…]
The results of the SDPED model with lengths 5 and 3 are listed below
SDPED_$\_5$(Ours) 0.660 0.672 0.616 0.676 0.692 0.645 0.683 0.693 0.649
SDPED_$\_3$(Ours) 0.654 0.668 0.617 0.674 0.687 0.647 0.684 0.695 0.658
Table 2:   Results on UDED with an error tolerance of 0.004 (equivalent to 1 pixel along the diagonal for the lowest-resolution image): The table is structured as Table 1. Across all three experiments and benchmarks, our model consistently achieves the best performance. The improvements over the second-best results are also reported. On average, our model (SDPED_$\_7$) improves scores by approximately $5.2\%$ in ODS, $4.4\%$ in OIS, and $2.4\%$ in AP.
 Partitions UDED-$P_{1}$-18-11-E150 UDED-$P_{2}$-20-9-E150 UDED-$P_{3}$-13-14-E150
 Benchmarks ODS OIS AP ODS OIS AP ODS OIS AP
HED (2015) 0.714 0.732 0.687 0.709 0.733 0.673 0.639 0.675 0.590
RCF (2017) 0.737 0.775 0.713 0.746 0.763 0.724 0.637 0.703 0.609
RIND (2021) 0.416 0.425 0.071 0.375 0.381 0.046 0.716 0.747 0.678
BDCN (2022) 0.767 0.787 0.763 0.772 0.803 0.757 0.689 0.762 0.686
Dexi (2023) 0.771 0.784 0.766 0.767 0.793 0.779 0.701 0.750 0.731
SDPED_$\_7$(Ours) 0.812 0.828 0.792 0.807 0.830 0.791 0.757 0.798 0.748
[…]
The results of the SDPED model with lengths 5 and 3 are listed below
SDPED_$\_5$(Ours) 0.811 0.824 0.795 0.806 0.827 0.796 0.756 0.793 0.744
SDPED_$\_3$(Ours) 0.807 0.822 0.786 0.803 0.825 0.798 0.753 0.791 0.751
Figure 5:  Sample predictions on BRIND and UDED, all from partition $P_{1}$ and displayed without NMS. The results illustrate that our model generates more visually coherent and reliable predictions.
4.3 Experiment Results
The evaluation results on BRIND, UDED datasets are presented in Table 1 and 2, while results on MDBD and BIPED2 are exhibited in Table 5 and 6, in supplemental materials. Additionally, some predicted images are displayed in Fig 5. Our models achieve SOTA performance compared to previous approaches.
On BRIND, our model improves the score by approximately $2.7\%$ in ODS, $3.1\%$ in OIS, and $6.9\%$ in AP on average. On UDED, the improvements are approximately $5.2\%$ in ODS, $4.4\%$ in OIS, and $2.4\%$ in AP [^8] Notably, RIND exhibits inconsistent performance across different test cases. While some results ($P_{1}$ and $P_{2}$) are significantly subpar, the other ($P_{3}$) performs better. This instability is evident in our retraining experiments: when retrained three times on $P_{2}$, RIND consistently yielded low AP scores. Similarly, retraining it on $P_{1}$ three times resulted in two low-AP outcomes. In contrast, RIND was retrained on $P_{3}$ only once, where it performed comparatively better. These results suggest that RIND may lack stability across different scenarios.. On MDBD, considering error tolerances of 0.001 (1.5 pixels) and 0.003 (4.4 pixels), our model improves performance by about $4.8\%/2.5\%$ in ODS, $4.8\%/2.9\%$ in OIS, and $22.5\%/14.7\%$ in AP, respectively. On BIPED2, our model significantly outperforms previous ones in terms of AP (improving by approximately $11.8\%$ on average), and meanwhile achieves comparable results in ODS and OIS, with differences of less than $0.3\%$ on average—regardless of whether our model or previous ones obtained the best scores.
4.4 Ablation studies
We conduct ablation studies on BRIND and UDED to evaluate the effectiveness of the skipping dense structure and the multiple-layer fusion block. Table 3 demonstrates that both components contribute positively to the performance.
Table 3: Ablation study of SDPED_$\_7$ on BRIND and UDED: No skipping refers to removing all skipping connections in SDB (excluding residual ones), while Single fuse replaces the multiple-layer fusion block with a single $1\times 1$ convolution layer. No skipping with single fuse removes all skipping connections and applies a single $1\times 1$ convolution layer for fusion, both as above. The best results are highlighted in bold, demonstrating the effectiveness of the proposed structures.
Partitions BRIND-$P_{1}$ UDED-$P_{2}$ BIPED2-$P_{3}$
Benchmark ODS OIS AP ODS OIS AP ODS OIS AP
SDPED_$\_7$ 0.660 0.672 0.613 0.807 0.830 0.791 0.677 0.682 0.598
No skipping 0.654 0.664 0.589 0.776 0.803 0.772 0.670 0.676 0.577
Single fuse 0.657 0.669 0.609 0.798 0.820 0.786 0.664 0.669 0.568
No skipping with single fuse 0.649 0.660 0.581 0.786 0.812 0.751 0.666 0.670 0.565
5 Discussion
In this section, we discuss the effectiveness of the SDPED architecture, along with its limitations and potential directions for future work.
5.1 Reasons for the Effectiveness of SDPED
The performance improvements of SDPED models, especially in AP, can be attributed to several factors:
5.1.1 Avoiding Information Loss
Most previous models incorporate down-sampling techniques, such as max-pooling, to extract multi-scale features. However, these operations inevitably lead to information loss. For instance, a $2\times 2$ max-pooling operation merges information from four pixels, making it impossible to recover fine details, thereby increasing uncertainty and reducing precision. In contrast, SDPED models eliminate down-sampling altogether, preserving all original information and mitigating uncertainty.
Down-sampling has traditionally been employed in ED models based on the assumption that multi-scale information is crucial, as demonstrated in classical approaches [9]. However, in deep CNN architectures, the receptive field is already sufficiently large to capture multi-scale features, rendering explicit down-sampling redundant. Since down-sampling not only removes information but also increases uncertainty, eliminating it in SDPED enhances precision and improves overall model performance.
5.1.2 Enhancing Texture Differentiation
The performance of ED models largely depends on their ability to distinguish between textures and true edges. However, determining the most effective architectural design for this task remains challenging due to the inherent ambiguity in defining textures. Notably, the backbone of SDPED has demonstrated strong texture-handling capabilities in SR tasks. Given that both SR and ED require precise identification of texture regions, SDPED’s architecture is naturally well-suited for ED. [^9] In SR tasks, the objective is to identify texture regions and enhance them, whereas, in ED tasks, the goal is to detect texture regions and suppress them. A model that excels in SR is likely to be effective in ED, as both tasks rely on accurate texture localization.
5.1.3 Preserving the Strongest Responses
Essentially, the distinction between edges and textures is based on response intensity to specific criteria, such as gradients. However, textural regions can also produce high responses, leading to false positives. SDPED addresses this issue through the use of CSDB, which can be viewed as an averaging mechanism across multiple SDB responses. This design reinforces pixels consistently identified as edges across multiple layers while suppressing less reliable responses. If a pixel exhibits a high response in only a single SDB layer but low responses in others, its final response is attenuated due to the skipping connections. As a result, SDPED effectively preserves only the most consistent edge responses while minimizing the influence of textural noise.
5.1.4 Strengthening Feature Fusion
ED architectures can be broadly categorized into feature extraction and feature fusion stages. While previous models allocate tens of millions of parameters to multi-scale feature extraction, they often rely on a single $1\times 1$ convolution layer for feature fusion. This stark imbalance between extraction and fusion limits the full utilization of extracted features. To address this limitation, SDPED incorporates an extended feature fusion block with additional layers, enhancing the integration of information and improving ED performance.
5.2 Model Size of SDPED
As shown in Table 4, SDPED has the fewest parameters among all compared models. Its model size is less than half that of previous architectures, which is particularly noteworthy given that it achieves comparable or superior performance. This result suggests that the SDPED architecture is more parameter-efficient, utilizing model weights more effectively than prior models. In contrast, the significantly larger number of parameters in previous models may indicate redundancy or inefficiencies in their design.
[…]
Size (kB) Num. of Coe. (×$\times 10^{6}$)
HED (2015) 57499 14.716171
RCF (2017) 57850 14.803781
RIND (2021) 232593 59.388526
BDCN (2022) 63749 16.302120
Dexi (2023) 137775 35.215245
SDPED_$\_7$(Ours) 22470 5.728638
5.3 Effectiveness Across Different ED Datasets
One of the key strengths of SDPED is its robustness across multiple ED datasets. As highlighted in [39], previous models often demonstrate dataset-specific performance, excelling on certain datasets while underperforming on others. For instance, HED and RCF achieve the highest accuracy on BRIND but are outperformed on UDED and BIPED2. Similarly, BDCN performs best on UDED but lags on other datasets, while Dexi and RIND yield superior results on MDBD and BIPED2 but fail to maintain consistency across different benchmarks.
In contrast, SDPED consistently achieves SOTA performance across various datasets. It yields the best results in all benchmarks on BRIND, UDED, and MDBD and significantly outperforms previous models in AP on BIPED2 while maintaining competitive performance on other benchmarks. This demonstrates its adaptability across different data distributions. These results indicate that SDPED is not only highly precise but also versatile, making it a robust solution for ED tasks across diverse datasets.
5.4 Limitations and Future Work
Despite the effectiveness of our methods, there remain areas for further improvement.
5.4.1 Inference Speed and Memory Consumption
Since SDPED processes full-resolution images, it requires more GPU memory and may have slower inference speeds compared to previous models. In this study, our primary focus was to enhance ED performance rather than optimize computational efficiency. However, in practical applications, a trade-off between accuracy and computational cost is often necessary. Future research could explore acceleration strategies and memory-efficient implementations to enhance real-time applicability.
5.4.2 Feature Fusion Schemes
Although ablation studies have shown the effectiveness of the extended feature fusion block, merely adding extra convolution layers may not be the optimal strategy for further improving edge detection performance. Since feature fusion has been largely overlooked in ED tasks, exploring more advanced fusion schemes could yield additional performance gains.
5.4.3 Noiseless ED Data
A fundamental challenge in ED tasks is the lack of a formal definition for edges. If edges had a strict mathematical definition, traditional algorithms would have already solved ED problems. As a result, ED datasets heavily rely on human-labeled annotations, which inherently introduce noise.
As demonstrated in Section IV, model performance is often correlated with dataset quality. For instance, all models perform best on UDED, despite it having the fewest training samples, likely due to its superior annotation quality. This suggests that data quality may be a more significant bottleneck than model architecture in ED tasks.
While Section III introduces an initial method for obtaining noiseless data, further research is needed to develop larger, high-quality datasets with minimal noise.
6 Conclusion
In this paper, we proposed SDPED, a novel class of precise ED models, and introduced a data augmentation strategy to improve model performance on noiseless data. Our model achieves SOTA performance across multiple datasets and partitions, particularly demonstrating significant improvements in AP. The results confirm that SDPED is both more precise and more adaptable than previous models. Furthermore, SDPED achieves superior performance with fewer parameters, utilizing less than half the coefficients of previous SOTA models. This highlights the parameter efficiency and architectural effectiveness of SDPED. Finally, our data augmentation strategy provides the first approach for leveraging noiseless data in training ED models, opening new possibilities for future research in obtaining high-quality ED data.
7 Bibliographies